# DQ-DETR: DETR with Dynamic Query for Tiny Object Detection

### v1

@@ 5 Conclusion
In this paper, we have analyzed the reasons why the previous DETR-like methods unsuitable for aerial datasets and proposed a new end-to-end transformer detector DQ-DETR with categorical counting module, counting-guided feature enhancement, and dynamic query selection. Our DQ-DETR dynamically adjust the number of object queries used for detection to solve the imbalance number of instances between different aerial images. Also, we improve the positional information of queries, and hence make the decoder easier to localize the tiny object. DQ-DETR is the first DETR-like model focusing on tiny object detection and achieves 30.2% $\mathrm{AP}$, which is the state-of-the-art of AI-TOD-V2. The result shows that our proposed DQ-DETR improves the performance of detecting tiny object, outperforming all previous CNN-based detectors and DETR-like methods on AI-TOD-V2 dataset with ResNet50 as the backbone.
[…]
Supplementary Material
6 Quantitative Result
6.1 FP/FN Under Different Density Situation
@@
{tblr}
width = column3 = c, column4 = c, column2 = c, cell21 = r=2, cell41 = r=2, vline2 = 1-2,4, vline2 = 3,5, hline1-2 = -, hline4,6 = -, Method Situation LRP FP LRP FN
@@
6.2 Ablation of Categorical Counting Module
@@
{tblr}
width = columneven = c, column3 = c, column5 = c, vline2 = -, hline1-2,5 = -, Method $\mathrm{AP}$ $\mathrm{AP_{vt}}$ $\mathrm{AP_{t}}$ $\mathrm{AP_{s}}$ $\mathrm{AP_{m}}$
@@
{tblr}
width = columneven = c, column3 = c, vline2-4 = -, hline1-2,7-8 = -, #Objects in image Accuracy(%) @ 4cls Accuracy(%) @ 5cls #Sample
@@
6.3 Results on VisDrone
Besides AI-TOD-V2, we also conduct experiments on the VisDrone [27] dataset to demonstrate the effectiveness of our model DQ-DETR.
VisDrone This dataset includes 14,018 drone-shot images, with 6,471 images for the train set, 548 for the validation set, and 3,190 for the test set. There are 10 categories, and the image resolution is 2000 $\times$ 1500 pixels. Also, the images are diverse in a wide range of aspects, including objects (pedestrians, vehicles, bicycles, etc.) and density (sparse and crowded scenes), where the average number of objects per image is 40.7 with a standard deviation of 46.41.
{tblr}
width = colspec = Q[438]Q[131]Q[158]Q[158], columneven = c, column3 = c, vline2 = -, hline1-2,9 = -, Model $\mathrm{AP}$ $\mathrm{AP_{50}}$ $\mathrm{AP_{75}}$
Faster R-CNN [16] 21.4 40.7 19.9
Cascade R-CNN [1] 22.6 38.8 23.2
Yolov5 [5] 24.1 44.1 22.3
CEASC [4] 28.7 50.7 24.7
SDP [11] 30.2 52.5 28.4
DINO-DETR* [26] 35.8 58.3 36.8
DQ-DETR (Ours) 37.0 60.9 37.9

### v2

@@ 4.6 Visualization
Fig. 2 visualizes the detection results of Deformable-DETRcompared with our DQ-DETR under sparse and dense situations. Deformable- DETR applies a fixed number(K=300) of queries, suffering a low recall rate under the imbalanced situation. The fixed number of queries will cause low detection accuracy in aerial datasets since the number of objects may vary drastically in different images. Using a small K restricts the recall of the objects in dense images, leaving lots of instances undetected (FN), as shown in Fig. 2(a). Conversely, Fig. 2(c) shows using larger K in the sparse images introduces many underlying false positive samples (FP).
@@ 5 Conclusion
In this paper, we analyze that the fixed number and position of queries in previous DETR-like methods are unsuitable for aerial datasets and propose a new end-to-end transformer detector DQ-DETR with categorical counting module, counting-guided feature enhancement, and dynamic query selection. Our DQ-DETR dynamically adjusts the number of object queries used for detection to solve the imbalance number of instances between different aerial images. Also, we improve the positional information of queries, making the decoder easier to localize the tiny object. DQ-DETR is the first DETR-like model focusing on tiny object detection and achieves 30.2% $\mathrm{AP}$, which is the state-of-the-art of AI-TOD-V2. The result shows that our proposed DQ-DETR improves the performance of detecting tiny objects, outperforming all previous CNN-based detectors and DETR-like methods on the AI-TOD-V2 dataset with ResNet50 as the backbone.
6 Supplementary
6.1 Quantitative Result
6.1.1 FP/FN Under Different Density Situation
@@
Method Situation LRP FP LRP FN
@@
6.1.2 Ablation of Categorical Counting Module
@@
Method $\mathrm{AP}$ $\mathrm{AP_{vt}}$ $\mathrm{AP_{t}}$ $\mathrm{AP_{s}}$ $\mathrm{AP_{m}}$
@@
#Objects in image Accuracy(%) @ 4cls Accuracy(%) @ 5cls #Sample
@@
6.2 The categorical counting module (CCM) is dataset distribution-depend.
Since the characteristics of different datasets may vary a lot, it is vital to tailor our categorical counting module depending on the dataset property and determined how many queries should be used for object detection. In categorical counting module for AI-TOD-V2, we estimate the counting number $N$, i.e., the number of instances per image, by a classification head and categorize them into four levels, which are $N\leq 10$, $10<N\leq 100$, $100<N\leq 500$, and $N>500$. It is worth noting that the hyperparameters 10, 100, 500 is tailored for AI-TOD-V2. For other datasets, the hyperparameters used in the CCM can be pipelined into a logical process using the mean and variance of the objects per image in the dataset to avoid long-tail distribution. There is no need to hand-design the CCM for different datasets.
Algorithm 1 Categorical Counting Pseudo Code
1:function Categorical-Counting($features$)
2:     $mean\leftarrow\text{Mean}(dataset)$
3:     $var\leftarrow\text{Variance}(dataset)$
4:     $GT\leftarrow\text{Predict}(features)$
5:     $class1\leftarrow\{GT<mean-var\}$
6:     $class2\leftarrow\{mean-var\leq GT<mean\}$
7:     $class3\leftarrow\{mean\leq GT<mean+var\}$
8:     $class4\leftarrow\{GT\geq mean+var\}$
9:end function
6.3 Visualization